OpenStudy (anonymous):

Challenge : For any natural number n, prove the following inequality : \(\large \mathsf{ \cfrac{1}{n+1 } (1+\cfrac{1}{3} + \cfrac{1}{5} + ...+ \cfrac{1}{2n-1} ) \ge \cfrac{1}{n} (\cfrac{1}{2} + \cfrac{1}{4} + ... +\cfrac{1}{2n} ) }\)

OpenStudy (anonymous):

1 < 2^1, so the proposition is true for n = 1. Assume it's true for some natural number k > 0. Then k < 2^k. Also 1 < 2^1 < 2^k. Adding k < 2^k and 1 < 2^k gives k + 1 < 2^k + 2^k = 2* (2^k) = 2^(k+1).

OpenStudy (anonymous):

Yes , but how does that disprove the given inequality Jack ?

OpenStudy (anonymous):

@FutureMathProfessor , how does that prove the given inequality.

OpenStudy (anonymous):

We are talking about natural numbers here Jack. \(\cfrac{1}{x} \) can not be considered here unless (unless x = 1 as it becomes a natural number then)

OpenStudy (jack1):

ah, my bad i assumed that natural numbers meant whole intergers

OpenStudy (dan815):